Tech Staff Meeting Notes April 30, 2007

Current Top Priorities are:Station Maintenance. ET. (Also wish to get INV populated.)

There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about

This task list as of September 12, 2006. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.

The Fiscal Fish Fry will be friday, June 29, 2007


Station Maintenance, NMSZ
Jim provided a spreadsheet of station/channel problems for NMSZ stations. COKM, TIPT, GRAT, and RDGT are fixed. BLAR is down. All the others are specific channel problems. Seem to be having a run on L-28 sensors. Able to troubleshoot to the geophone itself, replace it and repair the problem, then find nothing wrong with the replaced geophone (do I have that right Jim?). Weird. Several stations suffer from degrading telemetry signal strength due to growing trees. Might need taller masts that these stations.

SLU BB Stations. Little Rock is done but suspect its not as level as it should be. SLU needs another site visit but if anyone is in the area, contact Wurth to see if we can do it (e.g. James is going to Camp Robinson shortly). Other stations are on hold for various reasons. UT Martin is on deck. It is awaiting completion of unrelated construction which should be done in a week. Poplar Bluff is awaiting finalized permission. Ditto for Mountain Grove. SIUC ain't gonna happen.

Lenox Agreement. We have new contact info. Mitch will see if a new MOU is desired.

Got this note from Jim:
We installed the Freewave link between LCHS and prison yesterday. The antenna's are vertically polarized.

Also, the landowner uncertainty at Steel (STLE) appears to be resolved. We'll be on building power.
Power at LCHS, Lake Co School (is that the correct station James?) is drawn from a section of the grid that gets shut down during non-school hours. Need to isolate gps onto its own breaker not tied to the main that gets shut down. Also waiting on an ip for this site.
Finalizing bids for monument installation at two new Lake County sites.


Station Maintenance

John was headed to SMNC and CCRT during the meeting. Notes from last time are:
SMNC frequencies were changed. There is an additional problem. What is it?
CCRT has a large dc offset (permanently in the lowest gain setting) which causes trace to appear flat unless data are visually magnified. Also causes premature saturation on one side. Unfortunately the park is closed until April.

TZTN needs a gate. GPS antenna done.
GFM has a cable problem. Will visit in the spring.
Have some Lexan and batteries to give John.
MGNC appears to be experiencing some kind of interference.
The MYNC.US saga continues. Meremonte suspects the VSAT gateway. He's sending the entire stack for swap. John will need to flash the DRU with the DW4020 firmware and replace the stack. This never happened. Mitch will investigate.
Next trip of Memphis crew to ET is the week of May 7. Visits to Van Hill, Flat Gap, and the Morristown node (alarm install, batteries, and new pc).

Network Completion

The map shows stations with yellow triangles being new ones.
SMNC4done (see note above)
BHT8move (eventually)

C2SC is down. CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. James brought back some nice pictures from CVTN to help figure out how to do it (e.g. the conduit is way too small). Bill will get this done.
NAIT is the only non-real-time station. Wish to get it online. Might be able to hit Richardson from there.

Routine Maintenance Kudzu, battery check, etc, etc.

Second Tier Priorities

I-40 Bridge.
Need an 8 inch extension for surface vault to get it above grade.
Contractor got the sand out but had another snag. They have a fix for the snag and will implement shortly. Digitizers are in house. After much griping from Mitch, it looks like the best solution is to telemeter to the weigh station. From there we get on the TDOT fiber network. We telemeter from the TDOT POP at Boswell St (aka perimeter mall) to Richardson. A pc will be installed at richardson (unless there is a viable switch solution to transfer from private TDOT to CERI subnet). From there we hop on the CERI microwave and suck the data into anssworm. Hopefully the project has sufficient funds for the extra telemetry link (about $4k) and pc (about $1.5k). Chris is arranging to have a fence installed (after any pad, gravel or whatever is installed). Cost is $1100 installed (must be real silver). We don't want to do it ourselves since it has to tie into the existing dilapidated fence.

FDL. Got the new laptops. These are the rules:
They will all be configured exactly the same. Old laptops will be obsolesced NOT retasked. Each of the new ones will have a primary purpose (though if they're all the same, they should be able to fill all other purposes in a pinch):

  1. Aftershock deployments
  2. Environmental Geophysics Class
  3. Sundry other field work (e.g. mag, grav, etc)
  4. Bench
Presumably 1 and 4 could be secondary for each other and 2 and 3 could be secondary for each other.
Horton's laptop has all the necessary software. Need to pry it from is icy grip and extract the software and configs to the new laptops.
Two or three 6td's aren't working. Not sure whether its a config problem, cable problem, or sensor problem. Bill is investigating and returning all the rest to deployable status.
Steve will acquire a lacie pocket drive. If these can be made to work the same as the gizmos from guralp, he'll get another. The first will be for Bill's exclusive use (i.e. don't even let those pesky scientists know it exists else they'll want to take it into the field). Done, the drive works well and is about 4 times faster. Steve will get two more.
More requests to check-out magnetometer. What about software?
Seems to be resistance to check-out/in forms/procedures. Once we have all this in place, Mitch will be the hammer. There shall be no exceptions (not even Mitch and tech staff). This shall be true whenever something comes off the shelf whether its going to another State for deployment or just to another building at CERI.
Greg has surge supressors for remaining 6TD gps antennae. Bill will install.
Geometrics multichannel download will break when NT domain server goes away. This can be solved using NETBUI from the new laptops. James knows how to do this and Steve is very familiar with NETBUI. They'll show Bill the procedure and make sure it works.

As Time Permits

3890 Power Chris is going to make a diagram for 3890. Good luck son.

Generators P-plant was over the other week and, to our surprise, ran the generator (the new one). Greg informed them that it runs automatically every Monday morning. Not really a good idea to run it under no-load anyway. Doesn't really solve the need for routine mainenance like oil and filter changes, etc. Mitch will try to find out what their plans are.

Inventory Jim has all the broadband sensors defined and is working on additional equipment (e.g. isis) and the connections. Mitch will install Bill's account in mysql and inv so he can start building stations. Mitch needs to sit down with Jim and figure out what we need to have ISTI help us with. Sid will be available after about mid-May to finish up anything we need his help to take care of. Apparently the downloaded docs from guralp for the cmg5td's have the old +-4g contants. Mitch has paper documentation with the correct +-2g constants.

Steiner Brain Dump Seminar Marathon week while John was in Memphis. Completed 12 sessions so far. Can fit 1 hr/dvd and each dvd is 4.5 GB. Have about 12.5 hours so far. Would be nice to get these online. How public should they be?


All hardware is in house to replace the node computers in the coming year. We'll upgrade to ew v7.1 at the same time. This means we'll be using scnl instead of just scn. Steve, Al, and Mitch are working on the details (headaches) of that. Glasworm is the converter and is operational with the new switches.
Note that the new MTTN node pc is the exception. We're not ready for it to be scnlized yet.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data. Recentered, data are fine now.

VBMS. Sensor, DAS, VSAT, etc are installed. An additional trip was needed (and scheduled for last week) to tie up loose ends. Not sure if its operational.

Brent Rosenblad from Mizzou is good to go. Need to get dates from him and find out if its ok for CERI folks to observe.